# Lagrangian chaos and unique ergodicity for stochastic primitive equations

### v1

@@ 1.3. Overview
This manuscript is organized as follows.
  • In Section 2, we introduce the preliminary concepts related to PEs. Specifically, we recall the standard reformulation of the PEs as an evolution equation for the horizontal velocity $v_{t}$ alone. Additionally, we define $H^{\sigma}$-solutions to the stochastic PEs (1.1) and discuss the probabilistic framework.
  • In Section 3, we present the central results of this paper. More specifically, in Subsection 3.1 and Section 3.2, we outline the findings on the unique ergodicity of PEs, Lagrangian chaos, and scalar advection. Finally, in Section 3.3, we provide an overview of the proofs.
  • In Section 4, we prove the integrability condition (1.6) for invariant measures of the stochastic PEs (1.1).
  • In Section 5, we prove Theorem 1.2 and presenting complementary global well-posedness results for the stochastic PEs (1.1) in high-order Sobolev spaces.
  • In Section 6, we obtain the strong Feller property for the projective processes.
  • In Section 7, we establish approximate controllability for the Jacobian and projective processes, as well as weak irreducibility for the projective processes.
@@
Let $\mathbb{T}={\mathbb{R}}/{\mathbb{Z}}$. An element in the domain $\mathcal{O}$ is usually denoted by $\mathbf{x}=(x,y,z)$ where $(x,y)\in\mathbb{T}^{2}_{x,y}$ and $z\in(0,1)$ denotes the horizontal and the vertical directions, respectively. Here, $\mathbb{T}^{2}_{x,y}=\mathbb{T}\times\mathbb{T}$ and the subscript ‘$x,y$’ serving solely to indicate the corresponding coordinates. A similar decomposition is only employed on $\mathbb{T}^{3}=\mathbb{T}^{2}_{x,y}\times\mathbb{T}_{z}$ and $\mathbb{T}^{2}_{x,y}=\mathbb{T}_{x}\times\mathbb{T}_{y}$. The above splitting also induced a decomposition of the standard operators $\nabla=(\partial_{x},\partial_{y},\partial_{z})$ into the horizontal and vertical components given by $\nabla_{x,y}=(\partial_{x},\partial_{y})$ and $\partial_{z}$, respectively. In particular, for a vector field $v:\mathcal{O}\to{\mathbb{R}}^{2}$ with components $v_{x}$ and $v_{y}$, the operator $\nabla_{x,y}\cdot$ denotes the horizontal divergence and the action on $v$ is defined as $\nabla_{x,y}\cdot v=\partial_{x}v_{x}+\partial_{y}v_{y}.$
@@ 2.1. Reformulation and function analytic setting
As well-known in the study of PEs, they can be conveniently reformulated in terms of the only unknown $v$. To see this, note that the boundary condition at $w_{t}=0$ on $\mathbb{T}_{x,y}^{2}\times\{0\}$ in (1.1), combined with the incompressibility condition $\nabla_{x,y}\cdot v_{t}+\partial_{z}w_{t}=0$ leads to
@@
Summarizing the above discussion the PEs (1.1) can be reformulated in terms of the horizontal component of the velocity field $v:{\mathbb{R}}_{+}\times\Omega\times\mathbb{T}^{3}\to{\mathbb{R}}^{2}$ and the surface pressure $p:{\mathbb{R}}_{+}\times\Omega\times\mathbb{T}_{x,y}^{2}\to{\mathbb{R}}$ as
@@
Set $\mathbb{Q}_{x,y}f\stackrel{{\scriptstyle{\rm}}}{{=}}\nabla_{x,y}\Psi_{f}$. One can readily check that $\mathbb{Q}_{x,y}\in\mathscr{L}(L^{2}(\mathbb{T}^{2}_{x,y};{\mathbb{R}}^{2}))$. The hydrostatic Helmholtz projection $\mathbb{P}:L^{2}(\mathbb{T}^{3};{\mathbb{R}}^{2})\to L^{2}(\mathbb{T}^{3};{\mathbb{R}}^{2})$ is defined as
@@
To see the equivalence between (2.6) and (2.3) note that $\Delta$ commutes with $\mathbb{P}$, $\mathbb{P}(Q\dot{W})=Q\dot{W}$ (see the assumptions in Subsection 2.2.2 below) and the incompressibility condition (2.2) is satisfied as $\int_{\mathbb{T}_{z}}\nabla_{x,y}\cdot v(\cdot,z)\,\mathrm{d}z=0$ and such condition is preserved by the flow induced via (2.6).
@@ 2.2. Probabilistic set-up and $H^{\sigma}$ solutions
+${\mathbb{Z}}^{3}_{+}\cup{\mathbb{Z}}^{3}_{-}={\mathbb{Z}}^{3}_{0}\quad\text{ and }\quad{\mathbb{Z}}^{3}_{+}\cap{\mathbb{Z}}^{3}_{-}=\varnothing,$
where ${\mathbb{Z}}^{3}_{-}\stackrel{{\scriptstyle{\rm def}}}{{=}}-{\mathbb{Z}}^{3}_{+}$ and ${\mathbb{Z}}^{3}_{0}\stackrel{{\scriptstyle{\rm def}}}{{=}}{\mathbb{Z}}^{3}\setminus\{0\}$. Similar to [16, 70], we define the following real Fourier basis on $\mathbb{T}^{3}$:
$e_{\mathbf{k}}(\mathbf{x})\stackrel{{\scriptstyle{\rm def}}}{{=}}\left\{\begin{aligned} &\cos(2\pi\mathbf{k}\cdot\mathbf{x})&\ \ \ &\mathbf{k}\in{\mathbb{Z}}^{3}_{+}\ \text{ and }\ \mathbf{x}\in\mathbb{T}^{3},\\ &\sin(2\pi\mathbf{k}\cdot\mathbf{x})&\ \ \ &\mathbf{k}\in{\mathbb{Z}}^{3}_{-}\ \text{ and }\ \mathbf{x}\in\mathbb{T}^{3}.\end{aligned}\right.$
Let $\mathcal{P}\stackrel{{\scriptstyle{\rm def}}}{{=}}\{\mathbf{k}\in{\mathbb{Z}}^{3}_{0}\,:\,k_{z}=0\}$ where $k_{z}$ denotes the third component of the vector $\mathbf{k}$, i.e., $\mathbf{k}=(k_{x},k_{y},k_{z})$. Correspondingly, we define the set of indexes
@@
(2.11) $\gamma_{(\mathbf{k},\ell)}\stackrel{{\scriptstyle{\rm def}}}{{=}}\left\{\begin{aligned} &a_{\ell}&\qquad\text{ if }\ &(\mathbf{k},\ell)\in({\mathbb{Z}}_{0}^{3}\setminus\mathcal{P})\times\{1,2\},\\ &\mathbf{k}^{\perp}/|\mathbf{k}|&\qquad\text{ if }\ &(\mathbf{k},\ell)\in\mathcal{P}\times\{1\},\end{aligned}\right.$
where $\mathbf{k}^{\perp}=(k_{y},-k_{x})^{\top}$ for $\mathbf{k}\in\mathcal{P}$ and $a_{1}=(1,0)^{\top}$, $a_{2}=(0,1)^{\top}$ denotes the standard basis of ${\mathbb{R}}^{2}$. Finally, let us note that
@@ 2.2.1. $H^{\sigma}$-solutions
We can now give the definition of $H^{\sigma}$-solutions to (1.1) where $\sigma\geq 1$.
@@ 2.2.2. Probabilistic set-up
All the results of the current work hold under Assumption 2.2. However, we can weaken it for unique ergodicity results; see Subsection 3.1 below.
Comparing the above assumption with the one used in [16] (see Assumptions 1 and 2 in Subsection 1.1.1 there), one can expect Lagrangian chaos for PEs to hold even if $q_{(\mathbf{k},\ell)}=0$ for $N_{0}\leq|\mathbf{k}|\leq N_{1}$ for $N_{0},N_{1}>1$ (i.e., mid-frequency noise degeneracy). For simplicity, we do not pursue this level of generality here. However, we expect that the methods presented here and the existing literature on SPDEs with finite-dimensional noise degeneracy [29, 67, 70] can be combined to extend the results of the current manuscript to the latter situation. However, let us mention that we will, anyhow, encounter finite dimensional noise degeneracy in the study of strong Feller for the projective processes; see Section 6.
@@
As Proposition 3.1 below shows, the choice of $\sigma$ determines the pathwise regularity of the solution to the stochastic PEs (1.1). More precisely, the condition $\sigma<\alpha-\frac{3}{2}$ ensures $Q\in\mathscr{L}_{2}(\mathbb{L}^{2}_{0}(\mathbb{T}^{3}),\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3}))$ and therefore the solution $(v_{t})_{t}$ to (1.1) takes values in $C([0,\infty);H^{\sigma}(\mathbb{T}^{3};{\mathbb{R}}^{2}))$. The latter combined with $\sigma>\frac{9}{2}$ and $H^{\sigma}(\mathbb{T}^{3})\subseteq C^{\eta}(\mathbb{T}^{3})$, where $\eta=\sigma-\tfrac{3}{2}>3$, ensures the well-definedness of the projective process in (3.20)-(3.21). Finally, $\sigma>\alpha-\frac{3}{2}$ guarantees the uniqueness of the invariant measure for the stochastic PEs in (1.1); see Theorem 3.3 and Example 3.4.
@@ 3. Statement of the main results and strategy
In this section, we state our main results, and it is organized as follows:
@@ 3.1. Global well-posedness in $H^{\sigma}$ and unique ergodicity
We begin by stating the global well-posedness of the stochastic PEs (1.1) in high-order Sobolev spaces. Recall that a Markov process $(v_{t})_{t}$ is said to be Feller on $\mathbb{H}_{0}^{\sigma}(\mathbb{T}^{3})$ if the mapping
@@
where $(v_{t})_{t}$ is the solution to (1.1) with initial data $v\in\mathbb{H}_{0}^{\sigma}(\mathbb{T}^{3})$.
@@
Next, we present our results on the regularity and uniqueness of invariant measures for stochastic PEs (1.1). Let $\rho\geq 0$. A Borel probability measure $\mu$ on $\mathbb{H}_{0}^{\rho}(\mathbb{T}^{3})$ is called an invariant (or stationary) measure for (1.1) on $\mathbb{H}_{0}^{\rho}(\mathbb{T}^{3})$ if for all $t>0$ and $\phi\in C_{{\rm b}}(\mathbb{H}_{0}^{\rho}(\mathbb{T}^{3});{\mathbb{R}})$,
@@
Let $Q\in\mathscr{L}_{2}(\mathbb{L}_{0}^{2}(\mathbb{T}^{3}),\mathbb{H}_{0}^{\sigma}(\mathbb{T}^{3}))$ for some $\sigma\geq 2$, and fix $\rho\in[1,\sigma]$. Then there exists a strictly increasing continuous mapping $\Lambda:[0,\infty)\to[0,\infty)$ such that $\lim_{R\to\infty}\Lambda(R)=\infty$ such that for all invariant measures $\mu$ of the stochastic PEs (1.1) on $\mathbb{H}^{\rho}(\mathbb{T}^{3})$ satisfy, for all $\sigma^{\prime}<\sigma+1$,
(3.2) $\int_{\mathbb{H}^{\rho}_{0}(\mathbb{T}^{3})}\Lambda(\|v\|_{\mathbb{H}^{\sigma^{\prime}}(\mathbb{T}^{3})})\,\mathrm{d}\mu(v)\leq C(\|Q\|_{\mathscr{L}_{2}(\mathbb{L}_{0}^{2}(\mathbb{T}^{3}),\mathbb{H}_{0}^{\sigma}(\mathbb{T}^{3}))}).$
In particular, all invariant measures of the stochastic PEs (1.1) on $\mathbb{H}^{\rho}_{0}(\mathbb{T}^{3})$ are concentrated on $\bigcap_{\sigma^{\prime}<1+\sigma}\,\mathbb{H}^{\sigma^{\prime}}(\mathbb{T}^{3})$.
By parabolic regularity, the above is optimal. In its proof, we obtain (3.2) with $\Lambda(R)=\log^{[N]}(1+R)$ where $N\geq 1$ depends only on $\sigma$ and $\log^{[N]}(1+R)$ denotes the $N$-time composition of the mapping $R\mapsto\log(1+R)$. In case $\sigma$ is close to $2$, the previous bound is weaker than the one in [37, Theorems 1.7 and 1.9]. We do not expect this to be optimal, but it will be sufficient for our purposes.
@@
The previous answer positively to a question formulated by Glatt-Holtz, Kukavica, Vicol and Ziane below [37, Theorem 1.9].
Next, we turn our attention to uniqueness. As mentioned before Theorem 1.2, the following is the first uniqueness result for invariant measure for stochastic PEs. Recall that $v\in\mathrm{supp}\,\mu$ if $\mu(B_{\mathbb{H}^{\rho}_{0}(\mathbb{T}^{3})}(v,\delta))>0$ for all $\delta>0$.
@@
Suppose that $Q\in\mathscr{L}_{2}(\mathbb{L}_{0}^{2}(\mathbb{T}^{3}),\mathbb{H}_{0}^{\sigma}(\mathbb{T}^{3}))$ for some $\sigma\geq 2$. Moreover, assume that $Q$ has dense range and there exists $\gamma<\sigma+2$ for which the following nondegeneracy condition holds

### v2

@@ 1.3. Overview
This manuscript is organised as follows.
  • Section 2 introduces preliminary concepts for PEs, including the reformulation of PEs as an evolution equation for the horizontal velocity $v_{t}$ and the definition of $H^{\sigma}$-solutions to the stochastic PEs (1.1), along with the probabilistic framework.
  • Section 3 presents the main results of this paper: Unique ergodicity, Lagrangian chaos, and scalar advection by stochastic PEs in Subsection 3.1 and Section 3.2, with an overview of the proofs in Section 3.3.
  • Section 4 proves the integrability condition (1.6) for invariant measures of the stochastic PEs (1.1).
  • Section 5 proves Theorem 1.2 and complementary global well-posedness results for the stochastic PEs in high-order Sobolev spaces.
  • Section 6 establishes the strong Feller property for the projective processes.
  • Section 7 proves approximate controllability for the Jacobian and projective processes, along with weak irreducibility for the projective processes.
  • Appendix A provides complementary results on Markov semigroups.
@@
An element in the domain $\mathcal{O}$ is denoted by $\mathbf{x}=(x,y,z)$ where $(x,y)\in\mathbb{T}^{2}_{x,y}$ and $z\in(0,1)$ denotes the horizontal and the vertical directions, respectively. For $\mathbf{x}=(x,y,z)\in\mathcal{O}$, we denote by $\mathrm{dist}_{\partial\mathcal{O}}(\mathbf{x})$ its distance from the boundary:
$\mathrm{dist}_{\partial\mathcal{O}}(\mathbf{x})\stackrel{{\scriptstyle{\rm def}}}{{=}}z\wedge(1-z).$
In the above, $\mathbb{T}={\mathbb{R}}/{\mathbb{Z}}$ and $\mathbb{T}^{2}_{x,y}=\mathbb{T}\times\mathbb{T}$ and the subscript ‘$x,y$’ serving solely to indicate the corresponding coordinates. A similar decomposition is only employed on $\mathbb{T}^{3}=\mathbb{T}^{2}_{x,y}\times\mathbb{T}_{z}$ and $\mathbb{T}^{2}_{x,y}=\mathbb{T}_{x}\times\mathbb{T}_{y}$. The above splitting also induced a decomposition of the standard operators $\nabla=(\partial_{x},\partial_{y},\partial_{z})$ into the horizontal and vertical components given by $\nabla_{x,y}=(\partial_{x},\partial_{y})$ and $\partial_{z}$, respectively. In particular, for a vector field $v=(v_{x},v_{y})$ on either $\mathcal{O}$ or $\mathbb{T}^{3}$, the operator $\nabla_{x,y}\cdot$ denotes the horizontal divergence and the action on $v$ is defined as $\nabla_{x,y}\cdot v=\partial_{x}v_{x}+\partial_{y}v_{y}.$
@@ 2.1. Reformulation and function analytic setting
As is well-known in the study of PEs, they can be conveniently reformulated in terms of the only unknown $v$. To see this, note that the boundary condition at $w_{t}=0$ on $\mathbb{T}_{x,y}^{2}\times\{0\}$ in (1.1), combined with the incompressibility condition $\nabla_{x,y}\cdot v_{t}+\partial_{z}w_{t}=0$ leads to
@@
Summarising the above discussion, the PEs (1.1) can be reformulated in terms of the horizontal component of the velocity field $v:{\mathbb{R}}_{+}\times\Omega\times\mathbb{T}^{3}\to{\mathbb{R}}^{2}$ and the surface pressure $p:{\mathbb{R}}_{+}\times\Omega\times\mathbb{T}_{x,y}^{2}\to{\mathbb{R}}$ as
@@
Set $\mathbb{Q}_{x,y}f\stackrel{{\scriptstyle{\rm def}}}{{=}}\nabla_{x,y}\Psi_{f}$. One can readily check that $\mathbb{Q}_{x,y}\in\mathscr{L}(L^{2}(\mathbb{T}^{2}_{x,y};{\mathbb{R}}^{2}))$. The hydrostatic Helmholtz projection $\mathbb{P}:L^{2}(\mathbb{T}^{3};{\mathbb{R}}^{2})\to L^{2}(\mathbb{T}^{3};{\mathbb{R}}^{2})$ is defined as
@@
To see the equivalence between (2.6) and (2.3) note that $\Delta$ commutes with $\mathbb{P}$, $\mathbb{P}(Q\dot{W})=Q\dot{W}$ (see the assumptions in Subsection 2.2.2 below) and the incompressibility condition (2.2) is satisfied as $\int_{\mathbb{T}_{z}}\nabla_{x,y}\cdot v(\cdot,z)\,\mathrm{d}z=0$ and such a condition is preserved by the flow induced via (2.6).
@@ 2.2. Probabilistic set-up and $H^{\sigma}$ solutions
+${\mathbb{Z}}^{3}_{+}\cup{\mathbb{Z}}^{3}_{-}={\mathbb{Z}}^{3}_{0}\qquad\text{ and }\qquad{\mathbb{Z}}^{3}_{+}\cap{\mathbb{Z}}^{3}_{-}=\varnothing,$
where ${\mathbb{Z}}^{3}_{-}\stackrel{{\scriptstyle{\rm def}}}{{=}}-{\mathbb{Z}}^{3}_{+}$ and ${\mathbb{Z}}^{3}_{0}\stackrel{{\scriptstyle{\rm def}}}{{=}}{\mathbb{Z}}^{3}\setminus\{0\}$. Similar to [16, 71], we define the following real Fourier basis on $\mathbb{T}^{3}$ (as above, $\mathbb{T}={\mathbb{R}}/{\mathbb{Z}}$):
$e_{\mathbf{k}}(\mathbf{x})\stackrel{{\scriptstyle{\rm def}}}{{=}}2\left\{\begin{aligned} &\cos(2\pi\mathbf{k}\cdot\mathbf{x})&\ \ \ &\mathbf{k}\in{\mathbb{Z}}^{3}_{+}\ \text{ and }\ \mathbf{x}\in\mathbb{T}^{3},\\ &\sin(2\pi\mathbf{k}\cdot\mathbf{x})&\ \ \ &\mathbf{k}\in{\mathbb{Z}}^{3}_{-}\ \text{ and }\ \mathbf{x}\in\mathbb{T}^{3}.\end{aligned}\right.$
Let $\mathcal{P}\stackrel{{\scriptstyle{\rm def}}}{{=}}\{\mathbf{k}\in{\mathbb{Z}}^{3}_{0}\,:\,k_{z}=0\}$ where $k_{z}$ denotes the third component of the vector $\mathbf{k}$, i.e., $\mathbf{k}=(k_{x},k_{y},k_{z})$. Correspondingly, we define the set of indices
@@
(2.11) $\gamma_{(\mathbf{k},\ell)}\stackrel{{\scriptstyle{\rm def}}}{{=}}\left\{\begin{aligned} &a_{\ell}&\qquad\text{ if }\ &(\mathbf{k},\ell)\in({\mathbb{Z}}_{0}^{3}\setminus\mathcal{P})\times\{1,2\},\\ &\mathrm{sign}(\mathbf{k})\mathbf{k}^{\perp}/|\mathbf{k}|&\qquad\text{ if }\ &(\mathbf{k},\ell)\in\mathcal{P}\times\{1\},\end{aligned}\right.$
where $\mathbf{k}^{\perp}=(k_{y},-k_{x})^{\top}$ for $\mathbf{k}\in\mathcal{P}$ and $a_{1}=(1,0)^{\top}$, $a_{2}=(0,1)^{\top}$ denotes the standard basis of ${\mathbb{R}}^{2}$, and $\mathrm{sign}(\mathbf{k})=1$ if $\mathbf{k}\in{\mathbb{Z}}^{3}_{+}$ and $\mathrm{sign}(\mathbf{k})=-1$ otherwise. In particular, $\gamma_{(\mathbf{k},\ell)}=\gamma_{(-\mathbf{k},\ell)}$. One can check that
@@ 2.2.1. $H^{\sigma}$-solutions
We can now define $H^{\sigma}$-solutions to (1.1) with $\sigma\geq 1$.
@@ 2.2.2. Probabilistic set-up
All the results of the current work hold under Assumption 2.2. However, we can weaken it for unique ergodicity and regularity results; see Subsection 3.1 below.
Comparing the above assumption with the one used in [16] (see Assumptions 1 and 2 in Subsection 1.1.1 there), one can expect Lagrangian chaos for PEs to hold even if $q_{(\mathbf{k},\ell)}=0$ for $1<|\mathbf{k}|<N$ with $N>1$ (i.e., mid-frequency noise degeneracy). For simplicity, we do not pursue this level of generality here. However, we expect that the methods presented here and the existing literature on SPDEs with finite-dimensional noise degeneracy [29, 68, 71] can be combined to extend the results of the current manuscript to the latter situation. Let us mention that we will, anyhow, encounter finite-dimensional noise degeneracy in the study of strong Feller for the Lagrangian and projective processes; see Section 6.
@@
As Proposition 3.1 below shows, the choice of $\sigma$ determines the pathwise regularity of the solution to the stochastic PEs (1.1). More precisely, the condition $\sigma<\alpha-\frac{3}{2}$ ensures $Q\in\mathscr{L}_{2}(\mathbb{L}^{2}_{0}(\mathbb{T}^{3}),\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3}))$ and therefore the solution $(v_{t})_{t}$ to (1.1) takes values in $C([0,\infty);H^{\sigma}(\mathbb{T}^{3};{\mathbb{R}}^{2}))$. The latter combined with $\sigma>\frac{9}{2}$ and $H^{\sigma}(\mathbb{T}^{3})\subseteq C^{\eta}(\mathbb{T}^{3})$, where $\eta=\sigma-\tfrac{3}{2}>3$, ensures the well-definedness of the projective process in (3.20)-(3.21). Finally, $\sigma>\alpha-\frac{3}{2}$ guarantees the uniqueness of the invariant measure for on (1.1) on $\mathbb{H}^{\sigma}(\mathbb{T}^{3})$; see Theorem 3.3 and Example 3.4.
@@ 3. Statement of the main results and strategy
This section presents the main results and is organised as follows:
@@ 3.1. Global well-posedness in $H^{\sigma}$ and unique ergodicity
We begin by stating the global well-posedness of (1.1) in high-order Sobolev spaces. Recall that the Markov process $(v_{t})_{t}$ is said to be Feller on $\mathbb{H}_{0}^{\sigma}(\mathbb{T}^{3})$ if the mapping
@@
where the subscript in the expected value emphasises that $v_{0}=v$.
@@
Next, we present our results on the regularity and uniqueness of invariant measures for stochastic PEs (1.1). Let $\rho\geq 0$. A Borel probability measure $\mu$ on $\mathbb{H}_{0}^{\rho}(\mathbb{T}^{3})$ is called an invariant (or stationary) measure for (1.1) on $\mathbb{H}_{0}^{\rho}(\mathbb{T}^{3})$ if for all $t>0$ and for all continuous observable $\phi:\mathbb{H}_{0}^{\rho}(\mathbb{T}^{3})\to{\mathbb{R}}$, it holds that
@@
Let $Q\in\mathscr{L}_{2}(\mathbb{L}_{0}^{2}(\mathbb{T}^{3}),\mathbb{H}_{0}^{\sigma}(\mathbb{T}^{3}))$ for some $\sigma\geq 2$, and fix $\rho\in[1,\sigma]$. Then there exists a strictly increasing continuous mapping $\Phi:[0,\infty)\to[0,\infty)$ such that $\lim_{R\to\infty}\Phi(R)=\infty$ and for all invariant measures $\mu$ of the stochastic PEs (1.1) on $\mathbb{H}^{\rho}_{0}(\mathbb{T}^{3})$ satisfy, for all $\sigma^{\prime}<\sigma+1$,
(3.2) $\int_{\mathbb{H}^{\rho}_{0}(\mathbb{T}^{3})}\Phi(\|v\|_{\mathbb{H}^{\sigma^{\prime}}(\mathbb{T}^{3})})\,\mathrm{d}\mu(v)\leq C(\|Q\|_{\mathscr{L}_{2}(\mathbb{L}_{0}^{2}(\mathbb{T}^{3}),\mathbb{H}_{0}^{\sigma}(\mathbb{T}^{3}))}).$
In particular, all invariant measures of the stochastic PEs (1.1) on $\mathbb{H}^{\rho}_{0}(\mathbb{T}^{3})$ are concentrated on $\bigcap_{\sigma^{\prime}<1+\sigma}\,\mathbb{H}^{\sigma^{\prime}}_{0}(\mathbb{T}^{3})$.
By parabolic regularity, the above is optimal. We will obtain (3.2) with $\Phi(R)=\log^{[N]}(1+R)$ where $N\geq 1$ depends only on $\sigma$ and $\log^{[N]}(1+R)$ denotes the $N$-time composition of the mapping $R\mapsto\log(1+R)$. In case $\sigma$ is close to $2$, the previous bound is weaker than the one in [37, Theorems 1.7 and 1.9]. We do not expect this to be optimal, but it is sufficient for our purposes.
@@
The previous result provides an affirmative answer to a question posed by Glatt-Holtz, Kukavica, Vicol, and Ziane, stated below in [37, Theorem 1.9].
Next, we turn our attention to the uniqueness of invariant measures. As mentioned before Theorem 1.2, the following is the first uniqueness result for invariant measure for (1.1). Recall that $v\in\mathrm{supp}\,\mu$ if $\mu(B_{\mathbb{H}^{\rho}_{0}(\mathbb{T}^{3})}(v,\delta))>0$ for all $\delta>0$.
@@
Suppose that for some $\sigma\geq 2$ we have $Q\in\mathscr{L}_{2}(\mathbb{L}_{0}^{2}(\mathbb{T}^{3}),\mathbb{H}_{0}^{\sigma}(\mathbb{T}^{3}))$, the range of $Q$ is dense, and there exists $\gamma<\sigma+2$ such that